# Grounding Foundation Models through Federated Transfer Learning: A General Framework

### v10

@@ 6.3. Privacy and Backdoor Attacks
[…]
CrossLM (Deng et al., 2023) $\mathcal{W}^{C}$ $\mathcal{W}^{C}$ enhanced by $\mathcal{W}^{S}_{\text{FM}}$
FedLLM on Edge (Woisetschläger et al., 2023a) $\mathcal{W}^{C}$ Global model
@@
7. Co-optimize Foundation Models and Domain Models
The FTL-FM settings we discussed in Section 5 and Section 6 focus on adapting general knowledge of the FM to optimize DMs, and transferring domain-specific knowledge of DMs to optimize FM, respectively. In this section, we explore the setting 3⃝, which endeavors to mutually optimize FMs and DMs hosted by different parties (i.e., the server and clients). We summarize FTL-FM works that fall into this setting in Table 7, and in the following subsections, we will elaborate on these works.
7.1. Co-optimize FMs and DMs through model-level knowledge transfer
Optimizing both FMs and DMs helps establish a positive cycle to evolve FMs and DMs (DAMO, [n. d.]) continuously. For example, an FM server can first deliver the general knowledge and abilities of its FM to clients’ DMs, which are then trained on domain-specific data for downstream applications. Clients’ DMs, in turn, transfer industry-specific knowledge to enhance the server’s FMs. This loop can go beyond one training task and continues over time. Nevertheless, mutually enhancing FMs of the server and DMs of clients is rarely exploited in FL literature. A more commonly studied scenario involves clients collaboratively fine-tuning their domain models that are initialized with FMs. In this section, we consider both two scenarios:
  1. (1)
    The server owns an FM and clients own DMs. Clients’ DMs can optionally be initialized with FMs. The objective of this scenario is to optimize the server’s FM as well as the clients’ DMs. Consequently, the server ends up with an FM augmented with domain-specific knowledge, and each client with a DM enhanced with general knowledge of the FM.
  2. (2)
    The server serves as an aggregation function while clients’ DMs are initialized with or assisted by FMs. The objective of this scenario is to fine-tune to train clients’ DMs through federated learning. Each client ends up with a DM enhanced with the domain knowledge of all clients.
For the first scenario, FATE-LLM (Fan et al., 2023) proposed a FedCoLLM (Federated Co-tuning LLM) framework (Fan et al., 2023) to co-optimize the server’s FM and clients’ DMs simultaneously through mutual distillation. Figure 12(a) illustrates the workflow of FedCoLLM. In FedCoLLM, clients’ DMs are initialized with an off-the-shelf FM (i.e., LLaMA-7B (Touvron et al., 2023)). In each communication round, clients fine-tune their DMs using local data and then send them to the server with secure aggregation (Bonawitz et al., 2016). The server, in turn, distills the knowledge mutually between its hosted FM (i.e., LLaMa-65B) and the global DM aggregated from clients’ local DMs to co-optimize both models. Subsequently, the server dispatches the global DM to all clients for further training.
Deng et al. (2023) proposed CrossLM that optimizes the server’s FM and clients’ DMs through synthetic data that are generated by the server’s FM and validated by clients’ DMs, as illustrated in Figure 12(b). Specifically, in each communication round, clients train their DMs using local data and then send them to the server asynchronously. The server, in turn, generates synthetic data and leverages received DMs to validate the quality of synthetic data. Subsequently, the server utilizes validation feedback from DMs and the validated synthetic data to fine-tune the server’s FM. Next, the server uses the validated synthetic data to fine-tune each client’s DM and dispatches the fine-tuned DMs to corresponding clients for further training. This concerted effort is aimed at co-optimizing both the FM and DMs. FedCoLLM and CrossLM adhere to the objective formulated in Eq.(4) with $K>1$, $\sum^{K}_{k=1}p^{C_{k}}=1$, and $p^{S}=0$. FedCoLLM has $\eta>0$ whereas CrossLM has $\eta=0$. The specific form of the loss $\ell^{C_{k}}(\mathcal{W}^{S}_{\text{FM}},\mathcal{W}^{C_{k}};d)$ for co-optimizing the server’s $\mathcal{W}^{S}_{\text{FM}}$ and a client $k$’s $\mathcal{W}^{C_{k}}$ is formulated in Eq.(3.3).
Figure 12. Illustration of co-optimizing server’s FM and clients’ FM through transferring model-level knowledge among participating parties (i.e., server and client). (a) illustrates the federated co-optimization of clients’ full DMs and the server’s full FM. (b) illustrates federated parameter-efficient co-optimization of clients’ DMs and the server’s FM when the clients initialize their DMs with certain off-the-shelf FMs.
In addition to the model transfer method proposed by FedCoLLM, the representation transfer method proposed by CreamFL (Yu et al., 2023b) (illustrated in Figure 11(a)) also has the potential to achieve co-optimization since representations can convey general knowledge and contextual information of a model and thus can help improve the generalization capability of both FMs and DMs.
For the second scenario, FedNLP (Lin et al., 2022) provided a benchmarking framework for evaluating federated learning algorithms (e.g., FedAvg, FedProx, and FedOPT) on common NLP tasks with FMs (e.g., BART). Woisetschläger et al. (2023b) provided a micro-level hardware benchmark for measuring energy, computational, and communication efficiency when fine-tuning the FLAN-T5 model family, ranging from 80M to 3B parameters, using FL. Fine-tuning or training clients’ full FMs is computationally expensive, primarily because of the large sizes of FMs (Cai et al., 2023b). To alleviate this issue, various Parameter-Efficient Fine-Tuning (PEFT) methods have been proposed. As illustrated in Figure 13, these PEFT-based FL methods typically involve clients whose DMs are initialized with pre-trained FMs. During training, each client only uploads a PEFT module to the server for aggregation. The PEFT module can be an adapter, LoRA model, or prompt model that contributes to a small portion of the parameters of the original FM, thereby saving computational and communication overheads. These PEFT-based FL methods correspond to the objective formulated in Eq.(5) with $K>1$ and $\sum^{K}_{k=1}p^{C_{k}}=1$. The specific form of the loss $\ell^{C_{k}}(\mathcal{W}^{S}_{\text{FM}},\mathcal{W}^{C_{k}};d)$ for fine-tuning client $k$’s $\mathcal{W}^{C_{k}}$ with PEFT methods is formulated in Eq.(13).
@@ 7.1. Co-optimize FMs and DMs through model-level knowledge transfer
Figure 13. Federated parameter-efficient-finetuning of clients’ domain-specific FMs. Clients’ DMs are initialized with FMs and then learn from each other through federated fine-tuning of their PEFT modules (e.g., adapter, prompt model, and LoRA model). The server hosts no model and acts as an aggregation function.
@@
FedAdapter (Cai et al., 2023b) was proposed to improve the computational and communication efficiency of fine-tuning local FMs (e.g., BERT) resided in devices through PEFT (i.e., adapter). To achieve this, it dynamically adjusts the depth and width of adapters during training. FeS (Cai et al., 2023a) was proposed to address the problem of data scarcity in mobile NLP applications by utilizing pseudo-labeling. It employs a curriculum pacing mechanism to gradually speed up the pseudo-labeling speed, a representational filtering mechanism to pseudo-label samples that are most beneficial to learn from, and a co-planing mechanism to control the layer depth and capacity to be trained. Furthermore, FeS fine-tunes only bias parameters of FMs (e.g., RoBERTa-large) while keeping the other parameters frozen to improve training efficiency further. FwdLLM (Xu et al., 2023b) aims to enhance the efficiency of PEFT techniques. It observed that PEFT techniques mainly alleviate network bottlenecks but do not significantly improve training efficiency as they require backpropagation across the entire model to obtain gradients. To address this challenge, FwdLLM employs a backward propagation (BP)-free method to estimate gradients during the forward pass, thereby saving substantial memory and training time. To tackle system heterogeneity, Cho et al. (2023) proposed a heterogeneous LoRA (HLoRA) approach that can apply different rank LoRA modules to devices with heterogeneous and constrained computing resources by leveraging zero-padding and truncation for the aggregation. Woisetschläger et al. (2023a) conducted an in-depth study on hardware performance optimization for federated fine-tuning large language models at the edge. More specifically, It investigated the state-of-the-art deep learning accelerators as well as optimization levers for FL workloads on embedded hardware.
@@ 8. Foundation Model Inference
In literature, most of the privacy-preserving methods that aim to protect the user’s data privacy during inference focus on designing efficient Secure Multi-Party Computation (SMPC) methods (Li et al., 2023d; Dong et al., 2023; Liu and Liu, 2023; Zheng et al., 2023; Hou et al., 2023a; Hao et al., 2022; Ding et al., 2023). Given that SMPC comes at a high cost in terms of computational resources and communication, these methods primarily focus on optimizing the efficiency of FM architectures, predominantly Transformer-based neural networks, or SMPC protocols. We will comprehensively review these SMPC-based methods in Section 10. In a recent study, Tong et al. (2023) proposed InferDPT, which leverages DP-guaranteed text perturbation on the user’s prompts to prevent the adversary FM provider from inferring the user’s private data. Specifically, in the InferDPT framework, the user first randomly perturbs tokens of its prompt with semantically similar alternatives. Subsequently, the user sends the perturbed prompt to an LLM (i.e., GPT-4), which in turn generates a corresponding response and sends the response back to the user. Finally, the user employs a local pre-trained language model, such as Vicuna-7B, to generate the final response based on the original prompt and the response received from the LLM. DP-OPT (Hong et al., 2023) adopts Deep Language Network (DLN (Sordoni et al., 2023)) guided by a local FM to automatically optimize prompts. During the optimization, DP-OPT leverages a differentially private ensemble method to generate prompts to protect prompts’ privacy.  Staab et al. (2024) investigated leveraging text anonymization to protect the user’s data privacy, and it empirically demonstrated that anonymization is insufficient for protecting the user’s data privacy.
@@ 9.2.2. Adaptive Knowledge Transfer
Adaptive knowledge transfer (AKT) methods transfer knowledge from a teacher to a student through an adaptation procedure in which the transferred knowledge is selected based on the student’s data and task, as illustrated in Figure 16(c). This adaptability enables the student model to better capture the teacher model’s knowledge and perform knowledge transferring more efficiently.
@@ 10. Preserving Privacy
During federated transfer learning, participating parties transfer the knowledge of their private data and models from each other, aiming to ground FMs. Adversary parties in the loop can investigate transferred knowledge to infer the private information of other parties. Therefore, we need ways to protect transferred knowledge to mitigate the chance that adversaries can infer private information. We summarize data privacy protection methods adopted in FTL-FM works in Table 9.
@@
DP-SGD aims to train private deep learning models and utilizes privacy amplification to obtain a strong DP guarantee
(Abadi et al., 2016; Yu et al., 2022; Li et al., 2022d, 2023c; Shi et al., 2022b; Bu et al., 2022; Hou et al., 2023b)
DP-FTRL
DP variant of Follow-The-Regularized-Leader
(Kairouz et al., 2021; Wang et al., 2023c; Xu et al., 2023d)
Selective DP (SDP)
@@
SDP protects only the sensitive tokens defined by a policy function
(Shi et al., 2022a, b)
DP-based text privatization
@@
Protect the privacy of textual data based on DP
(Weggenmann et al., 2022; Li et al., 2023c; Zhou et al., 2023a; Igamberdiev et al., 2022; Igamberdiev and Habernal, 2023)
Cryptographic Protection Secure Aggregation (SA)
@@
SA is applied in FL to present the server from accessing model information of individual client
(WeBank, [n. d.])
Secure Multi-Party Computation (SMPC)
@@
Use SMPC to encrypt information transmitted between FL parties to protect data privacy
(Li et al., 2023d; Dong et al., 2023)
Homomorphic Encryption (HE)
@@
Use HE to encrypt information transmitted between FL parties to protect data privacy
(Chen et al., 2022a; Jin et al., 2023)
Hybrid
Hybrid of cryptographic protection methods
(Liu and Liu, 2023; Zheng et al., 2023; Hou et al., 2023a; Hao et al., 2022; Ding et al., 2023)
@@
Differential Privacy (DP) (Dwork et al., 2014) has become a widely used privacy protection mechanism because of its mathematical properties that facilitate the combination of multiple differential privacy mechanisms and the accumulation of privacy budgets. DP has several variants, including DP-SGD (Abadi et al., 2016), DP-FTRL (Kairouz et al., 2021), Selective-DP (SDP) (Shi et al., 2022a), and DP-Rewrite (Igamberdiev et al., 2022), that have been adopted to protect data privacy in FTL-FM works. DP-SGD (differentially private stochastic gradient descent) was proposed to train private deep learning models. It leverages privacy amplification by sampling or shuffling to achieve a strong formal DP guarantee. DP-SGD is not practical to be applied in a cross-device FL system because it requires the system to sample clients uniformly at random on each communication round to provide a formal DP guarantee. To address this obstacle,  Kairouz et al. (2021) proposed DP-FTRL (differentially private follow-the-regularized-leader) that can achieve privacy-utility-computation trade-offs comparable with DP-SGD but does not depend on privacy amplification. SDP (Selective DP) aims to improve the privacy-utility trade-off of DP-SGD when applied to natural language applications. It differs from DP-SGD in that SDP considers only partial dimensions of a training sample as sensitive, thereby saving privacy budgets to achieve the same model performance. DP-SGD, DP-FTRL, and SDP are typically applied to make neural network models DP-guaranteed, whereas DP-rewrite aims to make input text (e.g., textual documents) with DP guarantees by perturbing the original text representations. DP-rewrite techniques are convenient for protecting data-level knowledge transferred between parties in FTL-FM. While DP is a convenient tool for protecting data privacy, it often leads to noticeable utility deterioration, which weakens the motivation for adopting DP in utility-critical applications.

### v11

@@ 6.3. Privacy and Backdoor Attacks
7. Co-optimize Foundation Models and Domain Models
The FTL-FM settings we discussed in Section 5 and Section 6 focus on adapting general knowledge of the FM to optimize DMs, and transferring domain-specific knowledge of DMs to optimize FM, respectively. In this section, we explore the setting ③, which endeavors to mutually optimize FMs and DMs hosted by different parties (i.e., the server and clients).
Optimizing both FMs and DMs helps establish a positive cycle to evolve FMs and DMs (DAMO, [n. d.]) continuously. For example, an FM server can first deliver the general knowledge and abilities of its FM to clients’ DMs, which are then trained on domain-specific data for downstream applications. Clients’ DMs, in turn, transfer industry-specific knowledge to enhance the server’s FMs. This loop can go beyond one training task and continues over time. Nevertheless, mutually enhancing the server’s FM and clients’ DMs is rarely exploited in literature. A more commonly studied scenario involves clients collaboratively fine-tuning their domain models that are initialized with (or assisted by) FMs. We summarize the two scenarios as follows.
  1. (1)
    Co-optimize server’s FM and clients’ DMs. The server ends up with an FM augmented with domain-specific knowledge, and each client with a DM enhanced with general knowledge of the server’s FM. Note that, in this scenario, clients’ DM can be optionally initialized with FMs having sizes much smaller than the server’s FM.
  2. (2)
    Optimize clients’ domain-specific FMs. The server serves no FM but acts as a model aggregation function, while clients own DMs that are initialized with or assisted by FMs. The objective of this scenario is to optimize clients’ DMs through conventional horizontal federated learning. Each client ends up with a DM enhanced with the domain knowledge of all clients.
The FTL-FM methods proposed in the two scenarios almost focus on model-level knowledge transfer. We review FTL-FM methods of both scenarios in Section 7.1 and Section 7.2, respectively.
7.1. Co-optimize FMs and DMs through model-level knowledge transfer
FedCoLLM (Fan et al., 2023) and CrossLM (Deng et al., 2023) recently proposed in the literature fall into the first scenario. We summarize them in Table 7 and elucidate them as follows.
[…]
$\mathcal{D}^{C}$ $\mathcal{W}^{C}$ $\mathcal{D}^{S}$ $\mathcal{W}^{S}_{\text{FM}}$
[…]
Hetero-Model Transfer CrossLM (Deng et al., 2023) $\mathcal{W}^{C}$ $\mathcal{W}^{C}$ enhanced by $\mathcal{W}^{S}_{\text{FM}}$
FedCoLLM (Federated Co-tuning LLM) was proposed in FATE-LLM (Fan et al., 2023) (Federated Co-tuning LLM). It co-optimizes the server’s FM and clients’ DMs simultaneously through mutual knowledge distillation. Figure 12(a) illustrates the workflow of FedCoLLM. In FedCoLLM, clients’ DMs are initialized with an off-the-shelf FM (i.e., LLaMA-7B (Touvron et al., 2023)). In each communication round, clients fine-tune their DMs using local data and then send them to the server with secure aggregation (Bonawitz et al., 2016). The server, in turn, distills the knowledge mutually between its hosted FM (i.e., LLaMa-65B) and the global DM aggregated from clients’ local DMs. Subsequently, the server dispatches the global DM to all clients for further training. FedCoLLM adheres to the objective formulated in Eq.(4) with $K>1$, $\sum^{K}_{k=1}p^{C_{k}}=1$, $p^{S}=0$, and $\eta>0$. The specific form of the loss $\ell^{C_{k}}(\mathcal{W}^{S}_{\text{FM}},\mathcal{W}^{C_{k}};d)$ for co-optimizing the server’s $\mathcal{W}^{S}_{\text{FM}}$ and the global DM $\mathcal{W}^{G}$ shared by all clients is formulated in Eq.(13).
CrossLM was proposed by  Deng et al. (2023). It optimizes the server’s FM and clients’ DMs using synthetic data generated by the server’s FM and validated by clients’ DMs, as illustrated in Figure 12(b). Specifically, the DM on each client is randomly determined as BERT-Base or DistilBERT. In each communication round, clients train their DMs using local data and then send them to the server asynchronously. The server, in turn, generates synthetic data and leverages received DMs to validate the quality of synthetic data. Subsequently, the server utilizes validation feedback from DMs and the validated synthetic data to fine-tune the server’s FM. Next, the server uses the validated synthetic data to fine-tune each client’s DM and dispatches the fine-tuned DMs to corresponding clients for further training. This concerted effort is aimed at co-optimizing both the FM and DMs. CrossLM adhere to the objective formulated in Eq.(4) with $K>1$, $\sum^{K}_{k=1}p^{C_{k}}=1$, $p^{S}=1$, and $\eta=0$. The specific form of the tasks loss $\ell^{C_{k}}(\mathcal{W}^{S}_{\text{FM}},\mathcal{W}^{C_{k}};d)$ for co-optimizing the server’s $\mathcal{W}^{S}_{\text{FM}}$ and client $k$’s $\mathcal{W}^{C_{k}},k=1,\dots,K$ is formulated in Eq.(13).
Figure 12. Illustration of co-optimizing server’s FM and clients’ DMs through transferring model-level knowledge among participating parties (i.e., server and clients). (a) federated co-optimization is achieved through distilling knowledge between the server’s FM and the global DM aggregated from clients’ DMs. (b) federated co-optimization is achieved through fine-tuning the server’s FM and clients’ DMs using the synthetic data generated by the server’s FM and validated by clients’ DMs.
In addition to the model transfer methods, the representation transfer method proposed by CreamFL (Yu et al., 2023b) (illustrated in Figure 11(a)) also has the potential to achieve co-optimization since representations can convey general knowledge and contextual information of a model and thus can help improve the generalization capability of both FMs and DMs.
7.2. Optimize Domain-Specific FMs through model-level knowledge transfer
A considerable amount of FTL-FM research has been proposed in the second scenario, which essentially falls under the purview of horizontal federated learning (McMahan et al., 2017). FTL-FM works in this scenario achieve knowledge transfer typically through either homo-model transfer or hetero-model transfer. The homoe-model transfer requires all clients to initialize their DMs with the same FM and conduct knowledge transfer by sharing their full DMs with the server for aggregation. On the other hand, hetero-model transfer enables clients to initialize their DMs with FMs having different architectures and conduct knowledge transfer by sharing partial or proxies of their DMs. We summarize the two categories of FTL-FM works in Table 8 and elucidate them as follows.
@@ 7.2. Optimize Domain-Specific FMs through model-level knowledge transfer
FedNLP (Lin et al., 2022) Updates of DMs Global model
FedLLM-on-Edge (Woisetschläger et al., 2023) $\mathcal{W}^{C}$ Global model
@@
Optimizing clients’ domain-specific FMs through homo-model transfer corresponds to the objective formulated in Eq.(5) with $K>1$ and $\sum^{K}_{k=1}p^{C_{k}}=1$. The specific form of the task loss $\ell^{C_{k}}(\mathcal{W}^{G}_{\text{FM}};d)$ for fine-tuning the global model $\mathcal{W}^{G}_{\text{FM}}$ shared by all clients is formulated in Eq.(14). Current research of this category primarily revolves around benchmark design and privacy-preserving mechanisms. We will delve into the discussion of privacy-preserving mechanisms in Section 7.3 and examine the works related to benchmark design as follows.
FedNLP (Lin et al., 2022) provided a benchmarking framework for evaluating well-known federated learning algorithms (e.g., FedAvg, FedProx, and FedOPT) on FMs (e.g., BART) for common NLP tasks. Woisetschläger et al. (2023) conducted an in-depth study on hardware performance optimization for federated fine-tuning large language models on the edge (FedLLM-on-Edge). More specifically, It provided a micro-level hardware benchmark for measuring energy, computational, and communication efficiency when federated fine-tuning the FLAN-T5 model family, ranging from 80M to 3B parameters.
@@
Figure 13. Illustration of optimizing clients’ domain-specific FMs. In this scenario, clients’ DMs are initialized with FMs, while the server hosts no FM but acts as a model aggregation function. This scenario involves: (a) federated full-model fine-tuning of clients’ domain-specific FMs; (b) federated parameter-efficient-finetuning of PEFT modules (e.g., adapter, prompt model, and LoRA model) of clients’ domain-specific FMs.
One critical driving force behind the adoption of hetero-model transfer is the significant computational and communication costs associated with fine-tuning and sharing clients’ full domain-specific FMs. This is mainly due to the large sizes of FMs, as highlighted in the study by  Cai et al. (2023b). Current research on hetero-model transfer primarily focuses on the design and development of Parameter-Efficient Fine-Tuning (PEFT) methods, as exemplified in Table 8.
These PEFT-based FL methods, as illustrated in Figure 13, typically involve clients whose DMs are initialized with pre-trained FMs. During training, each client only uploads a PEFT module to the server for aggregation. The PEFT module can be an adapter, LoRA model, or prompt model that contributes to a small portion of the parameters of the original FM, thereby saving computational and communication overheads. These PEFT-based FL methods correspond to the objective formulated in Eq.(6) with $K>1$ and $\sum^{K}_{k=1}p^{C_{k}}=1$. The specific form of the task loss $\ell^{C_{k}}(\mathcal{W}^{S}_{\text{FM}},\mathcal{W}^{C_{k}};d)$ for fine-tuning client $k$’s $\mathcal{W}^{C_{k}}$ with PEFT methods is formulated in Eq.(15).
@@
FedAdapter (Cai et al., 2023b) was proposed to improve the computational and communication efficiency of fine-tuning local FMs (e.g., BERT) resided in devices through PEFT (i.e., adapter). To achieve this, it dynamically adjusts the depth and width of adapters during training. FeS (Cai et al., 2023a) was proposed to address the problem of data scarcity in mobile NLP applications by utilizing pseudo-labeling. It employs a curriculum pacing mechanism to gradually speed up the pseudo-labeling speed, a representational filtering mechanism to pseudo-label samples that are most beneficial to learn from, and a co-planing mechanism to control the layer depth and capacity to be trained. Furthermore, FeS fine-tunes only bias parameters of FMs (e.g., RoBERTa-large) while keeping the other parameters frozen to improve training efficiency further. FwdLLM (Xu et al., 2023b) aims to enhance the efficiency of PEFT techniques. It observed that PEFT techniques mainly alleviate network bottlenecks but do not significantly improve training efficiency as they require backpropagation across the entire model to obtain gradients. To address this challenge, FwdLLM employs a backward propagation (BP)-free method to estimate gradients during the forward pass, thereby saving substantial memory and training time. To tackle system heterogeneity, Cho et al. (2023) proposed a heterogeneous LoRA (HLoRA) approach that can apply different rank LoRA modules to devices with heterogeneous and constrained computing resources by leveraging zero-padding and truncation for the aggregation.
@@ 8. Foundation Model Inference
In literature, most of the privacy-preserving methods that aim to protect the user’s data privacy during inference focus on designing efficient Secure Multi-Party Computation (SMPC) methods (Li et al., 2023d; Dong et al., 2023; Liu and Liu, 2023; Zheng et al., 2023; Hou et al., 2023a; Hao et al., 2022; Ding et al., 2023). Given that SMPC comes at a high cost in terms of computational resources and communication, these methods primarily focus on optimizing the efficiency of FM architectures, predominantly Transformer-based neural networks, or SMPC protocols. We will comprehensively review these SMPC-based methods in Section 10. In a recent study, Tong et al. (2023) proposed InferDPT, which leverages DP-guaranteed text perturbation on the user’s prompts to prevent the adversary FM provider from inferring the user’s private data. Specifically, in the InferDPT framework, the user first randomly perturbs tokens of its prompt with semantically similar alternatives. Subsequently, the user sends the perturbed prompt to an LLM (i.e., GPT-4), which in turn generates a corresponding response and sends the response back to the user. Finally, the user employs a local pre-trained language model, such as Vicuna-7B, to generate the final response based on the original prompt and the response received from the LLM. DP-OPT (Hong et al., 2023) adopts Deep Language Network (DLN (Sordoni et al., 2023)) guided by a local FM to optimize prompts automatically. During the optimization, DP-OPT leverages a differentially private ensemble method to generate prompts to protect prompts’ privacy.  Staab et al. (2024) investigated leveraging text anonymization to protect the user’s data privacy, and it empirically demonstrated that anonymization is insufficient for protecting the user’s data privacy.
@@ 9.2.2. Adaptive Knowledge Transfer
Adaptive knowledge transfer (AKT) methods transfer knowledge from a teacher to a student through an adaptation procedure in which the transferred knowledge is selected based on the student’s data and task, as illustrated in Figure 16(c). This adaptability enables the student model to capture the teacher model’s knowledge better and perform knowledge transfer more efficiently.
@@ 10. Preserving Privacy
During federated transfer learning, participating parties transfer the knowledge of their private data and models from each other, aiming to ground FMs. Adversary parties in the loop can investigate transferred knowledge to infer other parties’ private information. Therefore, we need ways to protect transferred knowledge to mitigate the chance that adversaries can infer private information. We summarize data privacy protection methods adopted in FTL-FM works in Table 10.
@@
Differential Privacy (DP) DP-SGD DP-SGD aims to train private deep learning models and utilizes privacy amplification to obtain a strong DP guarantee (Abadi et al., 2016; Yu et al., 2022; Li et al., 2022d, 2023c; Shi et al., 2022b; Bu et al., 2022; Hou et al., 2023b)
DP-FTRL DP variant of Follow-The-Regularized-Leader (Kairouz et al., 2021; Wang et al., 2023c; Xu et al., 2023d)
@@
Selective DP (SDP) SDP protects only the sensitive tokens defined by a policy function (Shi et al., 2022a, b)
@@
DP-based text privatization Protect the privacy of textual data based on DP (Weggenmann et al., 2022; Li et al., 2023c; Zhou et al., 2023a; Igamberdiev et al., 2022; Igamberdiev and Habernal, 2023)
@@
Cryptographic Protection Secure Aggregation (SA) SA is applied in FL to present the server from accessing model information of individual client (WeBank, [n. d.])
@@
Secure Multi-Party Computation (SMPC) Use SMPC to encrypt information transmitted between FL parties to protect data privacy (Li et al., 2023d; Dong et al., 2023)
@@
Homomorphic Encryption (HE) Use HE to encrypt information transmitted between FL parties to protect data privacy (Chen et al., 2022a; Jin et al., 2023)
Hybrid Hybrid of cryptographic protection methods (Liu and Liu, 2023; Zheng et al., 2023; Hou et al., 2023a; Hao et al., 2022; Ding et al., 2023)
@@
Differential Privacy (DP) (Dwork et al., 2014) has become a widely used privacy protection mechanism because of its mathematical properties that facilitate the combination of multiple differential privacy mechanisms and the accumulation of privacy budgets. DP has several variants, including DP-SGD (Abadi et al., 2016), DP-FTRL (Kairouz et al., 2021), Selective-DP (SDP) (Shi et al., 2022a), and DP-Rewrite (Igamberdiev et al., 2022), that have been adopted to protect data privacy in FTL-FM works. DP-SGD (differentially private stochastic gradient descent) was proposed to train private deep learning models. It leverages privacy amplification by sampling or shuffling to achieve a strong DP guarantee. DP-SGD is not practical to be applied in a cross-device FL system because it requires the system to sample clients uniformly at random on each communication round to provide a formal DP guarantee. To address this obstacle,  Kairouz et al. (2021) proposed DP-FTRL (differentially private follow-the-regularized-leader) that can achieve privacy-utility-computation trade-offs comparable with DP-SGD but does not depend on privacy amplification. SDP (Selective DP) aims to improve the privacy-utility trade-off of DP-SGD when applied to natural language applications. It differs from DP-SGD in that SDP considers only partial dimensions of a training sample as sensitive, thereby saving privacy budgets to achieve the same model performance. DP-SGD, DP-FTRL, and SDP are typically applied to make neural network models DP-guaranteed, whereas DP-rewrite aims to make input text (e.g., textual documents) with DP guarantees by perturbing the original text representations. DP-rewrite techniques are convenient for protecting data-level knowledge transferred between parties in FTL-FM. While DP is a convenient tool for protecting data privacy, it often leads to noticeable utility deterioration, which weakens the motivation for adopting DP in utility-critical applications.